OpenStudy (anonymous):

A stone of mass 0.55 kilograms is released and falls to the ground. Measurements show that the stone has a kinetic energy of 9.8 joules at the point of impact with the ground. What is the stone's velocity when it hits the ground?

rvc (rvc):

You know the value of Kinetic Energy Also you know the mass of the stone K.E= 1/2 m v^2 therefore \(\rm v= \sqrt{\frac{2~K.E}{\Large m}}\)

rvc (rvc):

@thatone_girl :)